During a publicity campaign, a cycle shop gave away 5000 miniature cycles and bumper stickets. The cycles cost $0.21 each and the bumper stickers cost $0.14 each. The cycle stop spent $826 on the gifts. How many of each gift did they buy?
Answer by rfer(16322)

m+b=5000 b=5000-m
.21m+.14b=826
.21m+.14(5000-m)=826
.21m+700+.14m=826
.07m=126
m=1800
----------------
b=5000-1800
b=3200